The manager of an automobile plant would like to buy a certain machine to increase the output of his production line. He estimates the new machine will provide an added value of V = 160+40X^2 dollars per day, where X is a Poisson random variable with mean 0.96. On the other hand, the cost of operating the machine is 250$ per day. Is the manager making a good investment? Justify your answer by computing the expected profit/loss resulting from the new machine
Answer #1

E(V) = E(160 + 40X^2)

= 160 + 40 E(X^2)

= 0.96^2 + 0.96

= 1.8816

hence

E(V) = 160 + 40* 1.8816

= 235.264

since

E(V) < 250

the manager is not making good investment

Please rate